Convert 30 meter per second squared to km per hour squared


1
Expert's answer
2021-03-31T12:52:45-0400

Let's find how many of km/h^2 contains in one m/s^2:


"1\dfrac{m}{s^2} = 1\cdot \dfrac{1m}{1s^2} = 1\cdot \dfrac{10^{-3}km}{(1/3600)^2h^2}"

since "1m = 10^{-3}km", and "1s = 1\/3600h". Simplifying, obtain:


"1\dfrac{m}{s^2} = \dfrac{3600^2}{1000}\dfrac{km}{h^2} = 12960\dfrac{km}{h^2}"

Then 30 m/s^2 will be:


"30\dfrac{m}{s^2} = 30\cdot 12960\dfrac{km}{h^2} = 388800\dfrac{km}{h^2}"

Answer. "388800\dfrac{km}{h^2}".